# Every FLOP Counts: Scaling a 300B Mixture-of-Experts LING LLM without premium GPUs

### v1

@@ 1.2 Computing Environment for Model Training
The availability of computational resources is a critical determinant in the development of LLMs, particularly in the context of the increasingly popular MoE architecture. Recent State-of-the-art MoE models rely heavily on high-performance AI accelerators (e.g., H100 and H800) for training, yet the supply of such resources has remained constrained in recent years. Similar to the analysis on the imbalance of day-night inference load in the DeepSeek’s open-source files Team (2025a), in the commercial deployment of AI services, these high-performance resources are also in high demand during peak usage periods to ensure service quality. As a result, many LLM research organizations face persistent shortages of high-performance AI accelerators.
@@
  • Cross-cluster and cross-device compatibility. As described in the open-source project FlagScale Team (2025b), heterogeneous hardware environments often exhibit discrepancies in the implementation of low-level computational and communication operators and high-level distributed training frameworks. These challenges are particularly evident in training advanced architectures like MoE, where operators such as group_gemm, permute/unpermute, and all2all, and distributed strategies like expert parallelism may be missing or perform inconsistently across platforms. Ensuring training accuracy and portability requires: (1) collaboration with hardware vendors to standardize low-level operators, ensuring computational and communication consistency, (2) development of cross-platform compatibility layers to support seamless integration across distributed training frameworks, and (3) implementation of efficient debugging mechanisms for identifying and resolving issues in complex and heterogeneous environments.
@@ 1.3 Optimization for Model Training
  • Optimization of model architecture and training strategies. To enable efficient deployment on resource-constrained platforms, we adpot the following three strategies. (1) Model architecture optimization: Based on the comprehensive analysis of scaling laws for dense and MoE models, we can choose the best-matching architecture for the available computational resource. (2) Training framework optimization: For heterogeneous computing platforms, we integrate multiple training frameworks into a unified distributed deep learning framework, i.e., our open-source project, DLRover Team (2023). Additionally, to leverage the specific characteristics of various platforms, we develop a lightweight debugging tool, XPUTimer, which facilitates rapid and cost-effective task performance analysis while achieving a 90% reduction in memory usage. Furthermore, we implement a platform-agnostic asynchronous training strategy, namely EDiT (Elastic Distributed Training), which enhances the training efficiency, under various configurations, the training time can be reduced by up to 66.1%. (3) Storage optimization: Techniques such as device multi-tenancy and file system in user space (FUSE) are applied to achieve high performance and multi-cluster adaptability for large-scale training. Collaborative design of storage and training processes enhances I/O efficiency in MoE scenarios, reducing time overhead by 50%.
@@ 2.4 High-Efficiency Offline Inference Framework
The current mainstream inference frameworks are primarily designed for online inference, emphasizing increased throughput under specific latency constraints. Consequently, parallelization strategies predominantly include inter-node tensor parallelism (TP) and intra-node pipeline parallelism (PP) across multiple computational nodes. However, TP often incurs substantial communication overhead, particularly in computing systems that lack high-speed interconnects such as NVLINK. In such cases, communication overhead can account for more than half of the total execution time. To address these limitations and enhance throughput by reducing communication overhead, we propose an efficient offline inference framework named Flood Team (2025c), which adopts a fully pipeline-parallel (PP) architecture.
@@ 3.4.2 Long Context Pre-Training
During this phase of pre-training, the maximum input sequence length was extended to 16K tokens. This extension was achieved using Rotary Position Embedding (RoPE), with $\theta$ parameter adjusted from 10K to 600K to support longer sequences. Adjustments were also made to the training dataset to better align with the objectives of long-context processing. For the Ling-Plus model, the proportion of web-derived data was reduced, and additional long-form text data were incorporated to improve the model’s ability to process extended sequences. Similarly, for the Ling-Lite model, the amount of web-based data was scaled down, while the proportion of mathematical and coding-related corpora was increased. The learning rate schedule remained consistent with the prior training stage, and a total of 150B tokens were processed during this phase to enhance the model’s long-context processing capabilities. As a result, Figure 14 shows that Ling-Plus achieves nearly perfect performance in "Needle in A Haystack" testing across all context lengths up to 64K.
Figure 14: Needle in A Haystack Testing for Ling-Plus
@@ 3.4.3 Annealing
3.4.4 Skip loss spikes and retry mechanisms
Loss spikes are abrupt increases in training loss and are often caused by specific data and optimizer state combinations. Narrow, sharp spikes tend to have a minimal impact on performance, whereas wide, prolonged spikes can adversely affect both stability and model performance. During MoE training, such spikes may also result from hardware issues, such as malfunctioning accelerators or underperforming compute nodes. To mitigate these effects, we implemented a series of measures, including skip and retry mechanisms. When spikes occur for the first time, the affected update is skipped, the data is saved, and the training step is retried. If spikes persist upon retrying, we automatically reduce the learning rate during the affected step. Figure 15 intuitively illustrates the improvement in train loss achieved by the proposed strategy. This strategy has proven relatively effective in minimizing the impact of loss spikes compared to leaving them unaddressed.
@@ 3.4.4 Skip loss spikes and retry mechanisms
Figure 15: Comparison of train loss curves before and after applying skip spikes and retry mechanisms.
@@ 4.1.2 Data Redundancy
Specifically, we employed a text embedding model with demonstrated strong performance, as reported on the MTEB Leaderboard Team (2024a), to map problem prompts and responses into a vector space. Using these embeddings, we identified and removed instruction data with high cosine similarity, effectively reducing redundancy in the dataset. Our analysis revealed that removing approximately 10% to 20% of the most similar data had no adverse effect on the model’s core capabilities. This finding underscores the significant potential for deduplication within synthetic datasets to enhance data quality without compromising model performance.
@@ 4.3.1 Synthesis of Long-Context Instruction Data
To address the "lost in the middle" phenomenon Liu et al. (2024)—a common challenge in long-context tasks where key information located in the middle of lengthy documents is often overlooked—we constructed specialized datasets to improve model performance in these scenarios. Specifically, we created single-needle, multi-needle, and multi-hop retrieval datasets by inserting critical information into the middle of documents. These datasets were designed to enhance the model’s performance on needle-in-a-haystack tasks and related benchmarks requiring precise identification and retrieval of key information from extended contexts.
@@ 5.3.1 Evaluation Benchamarks
In addition to evaluating Ling models ability on various evaluation benchmarks, we also assess the models’ safety performance. Two evaluation datasets are constructed from the open-sourced data: 1) Arena Safety is constructed by randomly sampling 803 questions from a subset of lmsys-chat-1m Zheng et al. (2023) which is identified as risk by OpenAI moderation API Markov et al. (2023), and the responses are evaluated by Llama-Guard3 Team (2024b); 2) Cvalues Xu et al. (2023) uses 1711 multiple-choice questions to assess responsibility in a chinese context.

### v2

@@ 1.2 Computing Environment for Model Training
The availability of computational resources is a critical determinant in the development of LLMs, particularly in the context of the increasingly popular MoE architecture. Recent State-of-the-art MoE models rely heavily on high-performance AI accelerators (e.g., H100 and H800) for training, yet the supply of such resources has remained constrained in recent years. Similar to the analysis on the imbalance of day-night inference load in the DeepSeek’s open-source files DeepSeek (2025), in the commercial deployment of AI services, these high-performance resources are also in high demand during peak usage periods to ensure service quality. As a result, many LLM research organizations face persistent shortages of high-performance AI accelerators.
@@
  • Cross-cluster and cross-device compatibility. As described in the open-source project FlagScale FlagOpen (2025), heterogeneous hardware environments often exhibit discrepancies in the implementation of low-level computational and communication operators and high-level distributed training frameworks. These challenges are particularly evident in training advanced architectures like MoE, where operators such as group_gemm, permute/unpermute, and all2all, and distributed strategies like expert parallelism may be missing or perform inconsistently across platforms. Ensuring training accuracy and portability requires: (1) collaboration with hardware vendors to standardize low-level operators, ensuring computational and communication consistency, (2) development of cross-platform compatibility layers to support seamless integration across distributed training frameworks, and (3) implementation of efficient debugging mechanisms for identifying and resolving issues in complex and heterogeneous environments.
@@ 1.3 Optimization for Model Training
  • Optimization of model architecture and training strategies. To enable efficient deployment on resource-constrained platforms, we adpot the following three strategies. (1) Model architecture optimization: Based on the comprehensive analysis of scaling laws for dense and MoE models, we can choose the best-matching architecture for the available computational resource. (2) Training framework optimization: For heterogeneous computing platforms, we integrate multiple training frameworks into a unified distributed deep learning framework, i.e., our open-source project, DLRover DLRover (2023). Additionally, to leverage the specific characteristics of various platforms, we develop a lightweight debugging tool, XPUTimer, which facilitates rapid and cost-effective task performance analysis while achieving a 90% reduction in memory usage. Furthermore, we implement a platform-agnostic asynchronous training strategy, namely EDiT (Elastic Distributed Training), which enhances the training efficiency, under various configurations, the training time can be reduced by up to 66.1%. (3) Storage optimization: Techniques such as device multi-tenancy and file system in user space (FUSE) are applied to achieve high performance and multi-cluster adaptability for large-scale training. Collaborative design of storage and training processes enhances I/O efficiency in MoE scenarios, reducing time overhead by 50%.
@@ 2.4 High-Efficiency Offline Inference Framework
The current mainstream inference frameworks are primarily designed for online inference, emphasizing increased throughput under specific latency constraints. Consequently, parallelization strategies predominantly include inter-node tensor parallelism (TP) and intra-node pipeline parallelism (PP) across multiple computational nodes. However, TP often incurs substantial communication overhead, particularly in computing systems that lack high-speed interconnects such as NVLINK. In such cases, communication overhead can account for more than half of the total execution time. To address these limitations and enhance throughput by reducing communication overhead, we propose an efficient offline inference framework named Flood Flood (2025), which adopts a fully pipeline-parallel (PP) architecture.
@@ 3.4.2 Long Context Pre-Training
During this phase of pre-training, the maximum input sequence length was extended to 16K tokens. This extension was achieved using Rotary Position Embedding (RoPE), with $\theta$ parameter adjusted from 10K to 600K to support longer sequences. Adjustments were also made to the training dataset to better align with the objectives of long-context processing. For the Ling-Plus model, the proportion of web-derived data was reduced, and additional long-form text data were incorporated to improve the model’s ability to process extended sequences. Similarly, for the Ling-Lite model, the amount of web-based data was scaled down, while the proportion of mathematical and coding-related corpora was increased. The learning rate schedule remained consistent with the prior training stage, and a total of 150B tokens were processed during this phase to enhance the model’s long-context processing capabilities.
@@ 3.4.3 Annealing
3.4.4 Skip loss spikes and Sample retry mechanism
During pre-training, the phenomenon where the loss abruptly rises and then falls is referred to as loss spikes. These abrupt changes are typically triggered by specific interactions between the data and optimizer states. Loss spikes can be classified into two types: (1) narrow spikes, which last for only a few steps and have minimal impact on model performance, and (2) wide spikes, which persist across more steps and can significantly disrupt model stability, sometimes even causing benchmark evaluation results to approach random levels. Our research shows that it is difficult to completely eliminate loss spikes. To mitigate their effects, we have designed a series of strategies, including skip loss spikes and sample retry mechanism. When a loss spike is detected, the affected update is skipped, and the associated data is randomly re-injected into subsequent training batches. If the spike persists, we automatically reduce the learning rate during the affected step. This approach has proven effective in reducing the negative impact of loss spikes, enabling consistent improvements in benchmark metrics throughout the training process. Figure 14 intuitively illustrates the improvement in train loss achieved by the proposed strategy.
@@ 3.4.4 Skip loss spikes and Sample retry mechanism
Figure 14: Comparison of train loss curves before and after applying skip loss spikes and sample retry mechanism.
@@ 4.1.2 Data Redundancy
Specifically, we employed a text embedding model with demonstrated strong performance, as reported on the MTEB Leaderboard MTEB (2024), to map problem prompts and responses into a vector space. Using these embeddings, we identified and removed instruction data with high cosine similarity, effectively reducing redundancy in the dataset. Our analysis revealed that removing approximately 10% to 20% of the most similar data had no adverse effect on the model’s core capabilities. This finding underscores the significant potential for deduplication within synthetic datasets to enhance data quality without compromising model performance.
@@ 4.3.1 Synthesis of Long-Context Instruction Data
To address the "lost in the middle" phenomenon Liu et al. (2024)—a common challenge in long-context tasks where key information located in the middle of lengthy documents is often overlooked—we constructed specialized datasets to improve model performance in these scenarios. Specifically, we created single-needle, multi-needle, and multi-hop retrieval datasets by inserting critical information into the middle of documents. These datasets were designed to enhance the model’s performance on needle-in-a-haystack tasks and related benchmarks requiring precise identification and retrieval of key information from extended contexts. Figure 16 shows that Ling-Plus achieves nearly perfect performance in "Needle in A Haystack" testing across all context lengths up to 64K.
Figure 16: Needle in A Haystack Testing for Ling-Plus
@@ 5.3.1 Evaluation Benchamarks
In addition to evaluating Ling models ability on various evaluation benchmarks, we also assess the models’ safety performance. Two evaluation datasets are constructed from the open-sourced data: 1) Arena Safety is constructed by randomly sampling 803 questions from a subset of lmsys-chat-1m Zheng et al. (2023) which is identified as risk by OpenAI moderation API Markov et al. (2023), and the responses are evaluated by Llama-Guard3 PurpleLlama (2024); 2) Cvalues Xu et al. (2023) uses 1711 multiple-choice questions to assess responsibility in a chinese context.